OpenStudy (ttp):

Divide the quantity fifty-six times ten to the seventh divided by the quantity seven times ten to the twenty-first. Express the answer in scientific notation. eight times ten to the negative fourteenth eight times ten to the negative third eight times ten to the third eight times ten to the twenty-eighth

OpenStudy (anonymous):

fifty-six times ten to the seventh --------------------------- seven times ten to the twenty-first \[\frac{56 \times 10^{7}}{7 \times 10^{21}}\]ok, so if we divide the same base with an exponent, we have to ________ the exponents.

OpenStudy (ttp):

subtract the exponents?

OpenStudy (anonymous):

CORRECT so we do both a division 56/7, which is 8 and we subtract 21 from 10's exponent of 7

OpenStudy (anonymous):

7-21 = -14 8 x 10^-14

OpenStudy (ttp):

Lol THanks

OpenStudy (anonymous):

:)